OpenStudy (zarvox1):

4. What statement makes the open sentence 7x – 4 = 5x + 14 true? (Points : 1) x = 6 x = 9 x = 12 x = 8

OpenStudy (texaschic101):

7x - 4 = 5x + 14 --- subtract 5x from both sides 7x - 5x - 4 = 14 --- add 4 to both sides 7x - 5x = 14 + 4 -- combine like terms 2x = 18 -- divide both sides by 2 x = 9 <== this would make it true

OpenStudy (zarvox1):

thanks

OpenStudy (texaschic101):

no problem :)